OpenStudy (anonymous):

The length, l metres, of a field is 80m greater than the width. The perimeter is 600m. (i) Write the information in the form of an equation for 'l'. (ii) Solve the equation and so find the area of the field.

OpenStudy (anonymous):

let the width be x and length be x+80 perimeter=2(l+w) 600=2(x+80+x) 2x+80=600/2 2x=300-80 x=220/2 x=110 so,width=110 and length=110+80=190 area=l*w

OpenStudy (anonymous):

Fab answer, but i lost you half way through. Could you please label the answers as (i) and (ii)?

OpenStudy (anonymous):

where you lost ????

OpenStudy (anonymous):

After '600=2(x+80+x)'

OpenStudy (anonymous):

ok 600=2(x+80+x) 600/2=x+80+x 300=2x+80 300-80=2x 220=2x x=220/2 x=110

OpenStudy (anonymous):

Ohh I get it now :) Thanks...